Can the ratio of a parent's age to his or her child's age ever be exactly 10:9? explain?
Answer by jim_thompson5910(35256)

Let x = age of child and y = age of parent

Let's assume that the parent had the child at the age of 20 (or at least beyond that). So that would mean that the parent is 20 years older than the child. So that would mean that y = x+20

If the ratio of their ages is 10:9, then

y/x = 10/9

9y = 10x

9(x+20) = 10x

9x + 180 = 10x

180 = 10x - 9x

180 = x

x = 180

So if the parent was exactly 20 years older than the child, then the child would have to be 180 years old (the parent would be 200 years old).

This is definitely NOT likely. Well at least not likely in this current generation.